OpenStudy (1dfanatic):

Evaluate for: a = 2,k b = -1 and c = 1/3 (9c^2)^3

OpenStudy (anonymous):

(9(1/3)^2)^3 (1/3)^2=1/9 (1/9) times 9 is 1 1^3 = 1

OpenStudy (1dfanatic):

Thanks i was confused a bit.

OpenStudy (anonymous):

is that what you were looking for

OpenStudy (1dfanatic):

yeah :)

OpenStudy (anonymous):

step by step really simpify's it see ya next time